OpenStudy (anonymous):

solve each equation 1/x = 5/(x-4)

OpenStudy (anonymous):

1/x=5/(x-4) x-4=5x 4x=-4 x=-1

OpenStudy (anonymous):

how do you do the first step to get the x-4=5x that is what I'm lost at, I think I am making it harder than it is

OpenStudy (anonymous):



OpenStudy (anonymous):

oh ya duh I knew I was making it harder than it was

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw